Minutes of the 14th SOHO SWT Meeting

Leeuvenhorst Congres Centrum, Noordwijkerhout 27 February 1995 Contents

1 Action Items SWT 14 3 2 Agree Agenda and Actions Revision 3 3 Experiments Review 3 4 Payload & AIV Status (C. Berner) 4 5 SOHO Engineering Status (F. Felici) 5 6 Contingency planning (F. Felici) 5 7 SIM 2 (V. Domingo) 5 8 Data Rights (V. Domingo) 5 9 Cooperation with Ground-based Observatories (B. Fleck) 5 10 SOLAR PHYSICS Special (B. Fleck) 6 11 Science Working Team Executive Meeting 6 Annex 1: Agenda 7 Annex 2: List of Participants 8 Annex 3: Experiments Review 9

UVCS (D. Machi / J. Kohl) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CELIAS (D. Hovestadt) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 LASCO (G. Brueckner) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 EIT (J.-P. Delaboudiniere) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 SWAN (J.-L. Bertaux) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 SOI-MDI (P. Scherrer) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 COSTEP (H. Kunow) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 ERNE (J. Torsti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 SUMER (U. Schühle) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 CDS (R. Harrison) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 GOLF (A. Gabriel) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 VIRGO (C. Fröhlich) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

Annex 4: MCDR Task Force Report (L. Culhane) 168 Annex 5: Payload & AIV Status (C. Berner) 185 Annex 6: SOHO Engineering Status (F. Vandenbussche) 216 Annex 7: SIM 2 (V. Domingo) 232 Annex 8: Cooperation with Ground-based Observatories (B. Fleck) 241 Annex 9: SOLAR PHYSICS Special (B. Fleck) 244 Annex 10: SOHO Post-Launch PR Activties 246

2

1 Action Items SWT 14

Action 14.1 on Project to send information about SOHO model to PIs.

Action 14.2 on Project to send memo to PIs about accreditation at KSC.

Action 14.3 on PIs to provide ideas/“stories” for big SOHO media event. Input to PS before May 15.

Action 14.4 on PIs to comment on PR plan (draft 0.3 from 18 April 1995). Input to PS before May 15.

2 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2

Actions revision

Action 13.1 on PIs closed to submit NRT command time request for GSCT#2. Input to SOC before February 15.

Action 13.2 on SOC open to distribute coherent Early Ops Plan by 1 March.

Action 13.3 on PS closed to make Early Ops Plan an official document.

Action 13.4 on PIs closed to comment on list of activities involving S/C and other experiments. Input to SOC before 15 February.

Action 13.5 on PIs closed for CDS, UVCS, ERNE to provide their view on PR activities and a description of expected PR material from their experiment. Input to PS before 1 March.

Action 13.6 on PS closed to work out PR plan to be presented and discussed at next SWT.

Action 13.7 on PIs closed for SUMER to provide number of SOHO books they want to order. Input to B.Fleck before 1 May.

3 Experiments Review

In a special session held from 9 to 15 hours, the PI’s presented the status of their experiments to DSci, R.M. Bonnet (for hardcopies of the PI presentations see annex 3). The session was completed by L. Culhane who gave a summary of the MCDR Task Force report (annex 4).

D. Dale gave a summary of what happens now:

• UVCS

– agreed return to US – 29 May

– returned to KSC for re-integration – 24 July – planning available for rework

– ESA/NASA interaction: reporting/monitoring

• LASCO

– main problem: electronics unit exchange

– Project/PI review options and present for Level I decision 15/16 May

• EIT

– need of instrument removal & plans to be established by PI/Project based on 1. new filters

2. thermal regulation

3. door temperature

– meeting in ESTEC on May 10 (TBC)

– modifications to be implemented, if agreed, in Toulouse

• CELIAS

– MTOF and STOF flyable (100%) models to be supplied by PI at a date TBD in June (meeting setup for tomorrow 28 April)

– FM2 to be refurbished, retested and be available for possible reintegration in the last week of July (24 July).

• Other Instruments

– normal work

4 Payload & AIV Status (C. Berner)

see handout — annex 5

SOHO Experiments Review Experiments Acceptance Status Summaries (Rev. 1) was distributed.

PIs asked for details of planned activities at KSC. The Project agreed to send detailed information by week 19.^{1}

Note by P. Scherrer on control files updates required for EMC tests:

The first details of planned EMC tests were provided by M. Setzke on 23 March, along with a request for updates to control files. MDI responded 17 April, after significant work and verification with relatively minor changes. We were told on 20 April “there is no time in the schedule to debug control files or do extra tests”. We do not understand the approach of performing known insufficient tests just to get a test completed rather than in a manner to learn about the spacecraft and payload. The mission will suffer.

A. Poland brought up concerns expressed by PIs and FOT that planned test activties for GSCT#2 and 3 are not extensive enough to verify all experiment procedures. The Project replied that GSCTs are thought to test the ground segment of the operations chain and experiment procedures relevant for the commissioning of the experiments. A lively discussion developed. It was decided to discuss this issue more thoroughly at the upcoming SOWG on May 5.

1provided by Project on May 10

5 SOHO Engineering Status (F. Felici)

see handout — annex 6

C. Fröhlich questioned the validity of the thermal tests because of poorly known test environment. F. Felici disagreed and explained why.

6 Contingency planning (F. Felici)

F. Felici reported that Matra has been tasked to investigate major launch failure scenarios (under-/overburn of Centaur upper stage etc.) as well as some single point failures (solar array deployment, HGA, etc.) which may be recoverable.

7 SIM 2 (V. Domingo)

see handout - annex 7

8 Data Rights (V. Domingo)

V. Domingo presented the new SOHO Data Access Policy which should be acceptable to all parties involved. This policy is now included in the new issue of the SOP. It reads:

The intent of the SOHO data access policy is to provide data to as wide a community as possible and as soon as possible. From the beginning of the operational mission, the scientific community is welcomed and encouraged to participate in the analysis of the SOHO data in collaboration with the PI teams. The goal is to make fully callibrated data available for public use through ESA and NASA archives one year after reception by the PIs.

In addition to this general policy, the following rules apply:

• For each PI team the data distribution, data rights and publication policy is defined in their Science Book.

• All the PIs have the right to access the data of the other SOHO experimenters for planning purposes in the terms defined in section 3.4.2.1.

• For this purpose, each PI will make the data available according to a mutually agreed schedule.

• The SWT will establish rules to be applied to data obtained during agreed observation campaigns.

• Exchange of data acquired during internally coordinated SOHO observations will be regulated by the cooperating teams themselves.

9 Cooperation with Ground-based Observatories (B. Fleck)

see annex 8

This issue was extensively discussed after the PR meeting the day before the SWT on April 26 and was not re-discussed here due to lack of time. It will be discussed once more more thoroughly at the upcoming SPWG during SIM2 on May 2.

10 SOLAR PHYSICS Special (B. Fleck)

see annex 9

B. Fleck presented the status of the SOHO Special in Solar Physics. There are still 3 experiment descriptions missing: GOLF, EIT, and UVCS. To have the SOHO Special published in the September issue of Solar Physics, the complete and final camera-ready volume must be at the Kluwer office on June 1. Otherwise publication will slip to the November issue (October is out for other reasons). B.F. also needs number of hardcopy reprints for advance order (input from PI teams needed before May 15; cf. Action 13-7).

11 Science Working Team Executive Meeting

There are 4 action items from SWT-14 (see Sect. 1).

V. Domingo presented a summary of the PR meeting which took place the day before SWT-14 on May 26 at ESTEC (see annex 10). This meeting was attended by S. Vermeer, D. Nelson, M. Calabrese, W. Wagner, Ph. Lemaire, U. Schühle, R. Schwenn, G. Brueckner, D. Michels, C. Berner, J. Kohl, R. Harrison, J. Defise, H. Kunow, A. Poland, V. Domingo, B. Fleck, P. Scherrer, C. Fröhlich, R. Howard. S. Vermeer presented the SOHO Video and showed some typical examples of ESA/NASA video press releases. There it was agreed to form a SOHO PR Working Group (PRWG). Members of this PRWG will be S. Vermeer and D. Savage from ESA and NASA PR, the Project Manager, the Project Scientists, and several PIs or representatives. The following candidates were suggested: D. Gough and P. Scherrer for the helioseismology aspects, D. Michels or R. Schwenn, K. Wilhelm or U. Schühle, J. Kohl or P. Bochsler, and H. Kunow. The group should be formed during SIM2.^{2} At the SWT a discussion started about the PR involvement and priority rights of national funding agencies which payed for the instruments. No final conclusion was reached. This issue has to be discussed more thoroughly at one of the upcoming meetings during SIM2 and at the next SWT. There are two actions on the PIs: Action 14-3 to provide ideas/“stories” for the big SOHO media event which is planned to take place early after the end of the commissionung phase (input to PS before May 15), and Action 14-4 to comment on PR plan (draft 0.3 from 18 April 1995). Input to PS also before May 15.

No date fixed for next meeting.

2There and in subsequent discussions the following PI representatives in the PR Working Group were agreed upon: A. Gabriel, P. Scherrer, D. Michels, R. Schwenn, K. Wilhelm, J. Kohl, P. Bochsler, and H. Kunow

Annex 1: Agenda

7

Annex 2: List of Participants

8

Annex 3: Experiments Review

UVCS (D. Machi / J. Kohl)

CELIAS (D. Hovestadt)

37

LASCO (G. Brueckner)

51

EIT (J.-P. Delaboudiniere)

68

SWAN (J.-L. Bertaux)

70

SOI-MDI (P. Scherrer)

78

COSTEP (H. Kunow)

90

ERNE (J. Torsti)

99

SUMER (U. Schühle)

117

CDS (R. Harrison)

141

GOLF (A. Gabriel)

155

VIRGO (C. Fröhlich)

162

Annex 4: MCDR Task Force Report (L. Culhane)

168

Annex 5: Payload & AIV Status (C. Berner)

185

Annex 6: SOHO Engineering Status (F. Vandenbussche)

216

Annex 7: SIM 2 (V. Domingo)

232

Annex 8: Cooperation with Ground-based Observatories (B. Fleck)

241

Annex 9: SOLAR PHYSICS Special (B. Fleck)

244

Annex 10: SOHO Post-Launch PR Activties

246